OpenStudy (anonymous):

Passing through(7,5) with x-intercept -3. What is the slope intercept form?

OpenStudy (anonymous):

for point slope i get y-5=1/2(x-7) is this correct? and for some reason i keep getting the slope intercept wrong. What should that answer be?

OpenStudy (anonymous):

for point slope i get y-5=1/2(x-7) is this correct? ----> Yes

OpenStudy (anonymous):

By slope intercept , I assume you mean slope intercept equation. Just simplify the equation and express it as y=mx+c

OpenStudy (anonymous):

Any problem , reply back here @krissywatts

OpenStudy (anonymous):

thanks